# Sequential Change Detection with Differential Privacy

### v1

@@ Appendix B Proofs of main theorems
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{\frac{e^{(\frac{\lambda^{2}\Delta^{2}}{8}-I_{0}\lambda)i+\lambda\tilde{b}}}{1-4\Delta^{2}\lambda^{2}/\epsilon^{2}}}\leq 4(\frac{\Delta}{\epsilon})e^{\frac{1}{2}\lambda\tilde{b}}e^{-\frac{1}{2}(I_{0}\lambda-\frac{\lambda^{2}\Delta^{2}}{8})i}.$
Again taking $\lambda=\bar{\lambda}=\min\{4I_{0}/\Delta^{2},\epsilon/(2\sqrt{2}\Delta)\}$ guarantees $(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})>0$ and
Part III $\leq\frac{4\Delta}{\epsilon}e^{\frac{1}{2}\bar{\lambda}\tilde{b}}\sum_{i=m_{1}+1}^{\infty}e^{-\frac{1}{2}(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})i}\overset{(1)}{\leq}\frac{4\Delta}{\epsilon}e^{\frac{1}{2}\bar{\lambda}\tilde{b}}\frac{1}{\frac{1}{2}(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}e^{-\frac{1}{2}m_{1}(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}$
$\overset{(2)}{\leq}\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}e^{-\frac{1}{2}\tilde{b}\bar{\lambda}(1-\frac{{\bar{\lambda}}\Delta^{2}}{4I_{0}})}\leq\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})},$
@@
For Part II, note
Part II $\leq\sum_{i=m_{0}}^{m_{1}}(\mathbf{E}_{0}|Z_{i}|^{2})^{1/2}\sqrt{\mathbf{P}_{0}(\nu=i+1)}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\left(\sum_{i=m_{0}}^{m_{1}}\sqrt{\mathbf{P}_{0}(\nu=i+1)}\right)\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{(\sum_{i=m_{0}}^{m_{1}}1)(\sum_{i=m_{0}}^{m_{1}}\mathbf{P}_{0}(\nu=i+1))}$
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{m_{1}-m_{0}+1}\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{\frac{3\tilde{b}}{2I_{0}}+4}.$
For $b$ sufficiently large such that $b/(2I_{0})\geq 4$ we can further simplify the above expression as Part II $\leq 4\frac{\Delta}{\epsilon}\sqrt{\tilde{b}/I_{0}}$.
Step 4: Substituting everything into Eq. (25) we have when $b/(2I_{0})\geq 4$,
$\mathbf{E}_{0}[\nu]\leq\frac{\tilde{b}+4\frac{\Delta}{\sqrt{I_{0}}\epsilon}\sqrt{\tilde{b}}+\Delta+\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}+\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}+\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}}{I_{0}}=\frac{\tilde{b}+4\frac{\Delta}{\sqrt{I_{0}}\epsilon}\sqrt{\tilde{b}}+C}{I_{0}},$
where $C$ is a constant that depends only on $\epsilon$, $\Delta$, and $I_{0}$, but is independent of $b$.
Above we have shown the upper bound for $\mathbf{E}_{0}[\nu|W=w]$ when $w\geq 0$. Note that when $w<0$, we have $\mathbf{E}_{0}[\nu|W=w]\leq\mathbf{E}_{0}[\nu|W=|w|]$. Using $\sqrt{b+|w|}\leq\sqrt{b}+\sqrt{|w|}$, we have
$\mathbf{E}_{0}[\nu]$ $=\mathbf{E}_{W}\mathbf{E}_{0}[\nu|W=w]\leq\mathbf{E}_{W}[\frac{b+|W|+4\frac{\Delta}{\sqrt{I_{0}}\epsilon}(\sqrt{b}+\sqrt{|W|})+C}{I_{0}}]$
$=\frac{b}{I_{0}}+\frac{\mathbf{E}_{W}|W|}{I_{0}}+\frac{4\Delta}{I_{0}^{3/2}\epsilon}(\sqrt{b}+\mathbf{E}_{W}[\sqrt{|W|}])+C^{\prime}=\frac{b}{I_{0}}+\frac{4\Delta}{I_{0}^{3/2}\epsilon}\sqrt{b}+C^{\prime\prime},$
where $C^{\prime}$ and $C^{\prime\prime}$ are both constants independent of $b$. ∎
@@
$\mathbf{P}(\tilde{T}=t\mid X_{(1:t)\setminus k})$ $=\mathbf{P}(\tilde{T}=t\mid X_{(1:t)\setminus k},\mathcal{E}_{\delta})\mathbf{P}(\mathcal{E}_{\delta})+\mathbf{P}(\tilde{T}=t\mid X^{\prime}_{(1:t)\setminus k},\mathcal{E}_{\delta}^{c})\mathbf{P}(\mathcal{E}_{\delta}^{c})$
@@
This concludes the proof of Theorem 4.
@@
We first note that the proof of Lemma 2 remains valid, and thus the worst-case detection delay of $\tilde{T}$ is still attained when the changepoint occurs at $\tau=0$. Furthermore, the proof of Theorem 3 does not explicitly rely on the boundedness of the log-likelihood ratio $\ell(\cdot)$; it only invokes the boundedness of $\ell(X)$ through the use of Hoeffding’s inequality. As a result, the same proof applies in the unbounded case provided that $\ell(X)$ is sub-Gaussian under both the pre- and post-change distributions. Specifically, assume there exists a constant $\sigma^{2}>0$ such that $\mathbf{E}[e^{\lambda(\ell(X)-\mathbf{E}[\ell(X)])}]\leq e^{\frac{\lambda^{2}\sigma^{2}}{2}}$ for all $\lambda\in\mathbb{R}$. Under this condition, the same argument in the proof of Theorem 3 continues to hold with Laplace noise $\text{Lap}(2\Delta/\epsilon)$ replaced by $\text{Lap}(2A_{\delta}/\epsilon)$. The sub-Gaussian constant $\sigma^{2}$ appears only in the additive constant term $C$ in the resulting WADD bound. We omit the technical derivations, as they are essentially the same as the proof of Theorem 3.

### v2

@@ Appendix B Proofs of main theorems
$\leq 2\sqrt{2}(\frac{\Delta}{\epsilon})\sqrt{\frac{e^{(\frac{\lambda^{2}\Delta^{2}}{8}-I_{0}\lambda)i+\lambda\tilde{b}}}{1-4\Delta^{2}\lambda^{2}/\epsilon^{2}}}\leq 4(\frac{\Delta}{\epsilon})e^{\frac{1}{2}\lambda\tilde{b}}e^{-\frac{1}{2}(I_{0}\lambda-\frac{\lambda^{2}\Delta^{2}}{8})i},$
where the last inequality is due to $1-4\Delta^{2}\lambda^{2}/\epsilon^{2}\geq\frac{1}{2}$ when $\lambda\leq\frac{\epsilon}{2\sqrt{2}\Delta}$. Taking $\lambda=\bar{\lambda}=\min\{4I_{0}/\Delta^{2},\epsilon/(2\sqrt{2}\Delta)\}$ guarantees $(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})>0$ and thus,
Part II $\leq\frac{4\Delta}{\epsilon}e^{\frac{1}{2}\bar{\lambda}\tilde{b}}\sum_{i=m_{1}+1}^{\infty}e^{-\frac{1}{2}(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})i}$
$\overset{(1)}{\leq}\frac{4\Delta}{\epsilon}e^{\frac{1}{2}\bar{\lambda}\tilde{b}}\frac{1}{\frac{1}{2}(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}e^{-\frac{1}{2}m_{1}(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}$
$\overset{(2)}{\leq}\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})}e^{-\frac{1}{2}\tilde{b}\bar{\lambda}(1-\frac{{\bar{\lambda}}\Delta^{2}}{4I_{0}})}\leq\frac{8\Delta}{\epsilon(I_{0}\bar{\lambda}-\frac{{\bar{\lambda}}^{2}\Delta^{2}}{8})},$ (28)
@@
Step 4: Finally, we substitute the upper bound for $\mathbf{E}_{0}[|Z_{\nu-1}|]$ as derived in Step 3 into Eq. (25) to complete the proof. Substituting the upper bounds in (27) and (28) into Eq. (26) and (25), we have for any $\tilde{b}>0$,
[…]
$=\frac{\tilde{b}+{4\frac{\Delta}{\sqrt{I_{0}}\epsilon}}\sqrt{\tilde{b}}+C}{I_{0}},$
where $C$ is a constant that depends only on $\epsilon$, $\Delta$, and $I_{0}$, but is independent of $\tilde{b}$.
Above we have shown the upper bound for $\mathbf{E}_{0}[\nu(b+w)|W=w]$ for any $b\geq 0$ when $w\geq 0$. Note that when $w<0$, we have $\mathbf{E}_{0}[\nu(b+w)|W=w]\leq\mathbf{E}_{0}[\nu(b+|w|)|W=|w|]$. Using $\sqrt{b+|w|}\leq\sqrt{b}+\sqrt{|w|}$, we have
$\mathbf{E}_{0}[\tilde{T}(b)]$ $=\mathbf{E}_{W}\mathbf{E}_{0}[\nu(b+w)|W=w]$
$\leq\mathbf{E}_{W}\mathbf{E}_{0}[\nu(b+|w|)|W=w]$
$\leq\mathbf{E}_{W}[\frac{b+|W|+4\frac{\Delta}{\sqrt{I_{0}}\epsilon}(\sqrt{b}+\sqrt{|W|})+C}{I_{0}}]$
$=\frac{b}{I_{0}}+\frac{\mathbf{E}_{W}|W|}{I_{0}}+\frac{4\Delta}{I_{0}^{3/2}\epsilon}(\sqrt{b}+\mathbf{E}_{W}[\sqrt{|W|}])+C^{\prime}$
$=\frac{b}{I_{0}}+\frac{4\Delta}{I_{0}^{3/2}\epsilon}\sqrt{b}+C^{\prime\prime},$
where $C^{\prime}$ and $C^{\prime\prime}$ are both constants independent of $b$.
@@
$\mathbf{P}(\tilde{T}=t\mid X_{(1:t)\setminus k})$
$=\mathbf{P}(\tilde{T}=t\mid X_{(1:t)\setminus k},\mathcal{E}_{\delta})\mathbf{P}(\mathcal{E}_{\delta})$
$\hskip 10.0pt+\mathbf{P}(\tilde{T}=t\mid X^{\prime}_{(1:t)\setminus k},\mathcal{E}_{\delta}^{c})\mathbf{P}(\mathcal{E}_{\delta}^{c})$
@@
This concludes the proof of Theorem 4. ∎
@@
We first note that the proof of Lemma 2 remains valid, and thus the worst-case detection delay of $\tilde{T}$ is still attained when the changepoint occurs at $\tau=0$. Furthermore, the proof of Theorem 3 does not explicitly rely on the boundedness of the log-likelihood ratio $\ell(\cdot)$; it only invokes the boundedness of $\ell(X)$ through the use of Hoeffding’s inequality. As a result, the same proof applies in the unbounded case provided that $\ell(X)$ is sub-Gaussian under both the pre- and post-change distributions. Specifically, assume there exists a constant $\sigma^{2}>0$ such that $\mathbf{E}[e^{\lambda(\ell(X)-\mathbf{E}[\ell(X)])}]\leq e^{\frac{\lambda^{2}\sigma^{2}}{2}}$ for all $\lambda\in\mathbb{R}$. Under this condition, the same argument in the proof of Theorem 3 continues to hold with Laplace noise $\text{Lap}(2\Delta/\epsilon)$ replaced by $\text{Lap}(2A_{\delta}/\epsilon)$. The sub-Gaussian constant $\sigma^{2}$ appears only in the additive constant term $C$ in the resulting WADD bound. We omit the technical derivations, as they are essentially the same as the proof of Theorem 3. ∎